OpenStudy (anonymous):

Write the Riemann sum to find the area under the graph of the function f(x) = x^2 from x = 1 to x = 5: (I'm going to post what I wrote and I wanted to see if it was correct)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \sum_{i \rightarrow 1}^{n} (\frac{ i }{ n })^{2}(\frac{ 4 }{ n })\] (now let me just explain why I think it's this (though please tell me ahead of time if it's wrong))

OpenStudy (mathmale):

Fun problem. Are you planning to use "left end points" or "right end points?"

jimthompson5910 (jim_thompson5910):

it's close, but not quite there

OpenStudy (anonymous):

I wrote my answer based on the idea of using left end points, but I'm not completely sure that's what they wanted. Also I don't have the whole riemann sum and limits thing down.

OpenStudy (anonymous):

endpoints make no difference what does make a difference is that you are starting at \(1\)

jimthompson5910 (jim_thompson5910):

also, slight notation note, but see attached

jimthompson5910 (jim_thompson5910):

anyways, if you were starting at x = 0, then you'd have the correct answer

OpenStudy (mathmale):

If you're using a finite number of x values, then the use of left- or right-end-points does make a difference. However, if you intend to find the exact value of the area defined by the curve, then the choice makes no difference.

OpenStudy (anonymous):

Okay. Sorry if I'm being a bit slow.

OpenStudy (mathmale):

\[\lim_{x \rightarrow \infty} \sum_{i \rightarrow 1}^{n} (\frac{ i }{ n })^{2}(\frac{ 4 }{ n })\]

OpenStudy (mathmale):

should be \[\lim_{n \rightarrow \infty} \sum_{i \rightarrow 1}^{n} (\frac{ i }{ n })^{2}(\frac{ 4 }{ n })\]

OpenStudy (anonymous):

I'm trying to understand what I'm doing

OpenStudy (mathmale):

...where I've changed the 'x' to an 'n.'

OpenStudy (anonymous):

Oh yeah, that was just an accident

OpenStudy (mathmale):

Could you identify what's not yet clear for you? I'm sure it was just an accident.

OpenStudy (anonymous):

Okay, first- I know I'm supposed to use the limit, one because they don't ask for an approximation and two because this is what the lesson is about. I know the (4/n) part is correct because the width get's infinity smaller/ approaching zero, and that would give me the exact area. The middle part that's neither that nor the limit I sort of took from my notes but I'm not entirely sure how that functions.

OpenStudy (mathmale):

Suggestion: If you haven't already, set up a formula for the x values. It will look like\[x _{i}=a+i \frac{ b-a }{ n }\]

jimthompson5910 (jim_thompson5910):

Left Endpoint Riemman Sum \[\Large A = \lim_{n \to \infty}\left[\sum_{i=0}^{n-1}f(x_i)*\Delta x\right]\] Right Endpoint Riemman Sum \[\Large A = \lim_{n \to \infty}\left[\sum_{i=1}^{n}f(x_i)*\Delta x\right]\] If you don't include the limit as n goes to infinity, the two Riemman sums will be different. If you do let n go to infinity, then the two areas are equal. A: area under the curve from x = a to x = b i: index value n: number of rectangles \(\Large x_i = a + \Delta x*n\) \(\Large \Delta x = \frac{b-a}{n}\)

OpenStudy (mathmale):

Here your a=1 and your b=5.

jimthompson5910 (jim_thompson5910):

oops typo, I meant to say \(\Large x_i = a + \Delta x*i\)

OpenStudy (anonymous):

Oh okay, do you guys mind if I give it another go thanks a ton btw

OpenStudy (anonymous):

that really helps

OpenStudy (mathmale):

Of course. Thanks for your persistence and your willingness to try things on your own. Jim, many thx for your contributions here.

jimthompson5910 (jim_thompson5910):

np, you'll also use these summation identities https://www.math.ucdavis.edu/~kouba/CalcTwoDIRECTORY/summationdirectory/Summation.html mainly the first three in this case

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \sum_{i=1}^{n} (1 + \frac{ 4i }{ n })^{2}(\frac{ 4 }{ n })\]

OpenStudy (mathmale):

Nice job of expressing x_sub_1!

OpenStudy (anonymous):

(thanks so much, the majority of the stuff you guys gave me is going in my notes)

OpenStudy (mathmale):

Truly happy for you.

OpenStudy (tkhunny):

Work on your parentheses, though. ( \ d f r a c { i } { n } ) ^ { 2 } produces \((\dfrac{i}{n})^{2}\) \ l e f t ( \ d f r a c { i } { n } \ r i g h t ) ^ { 2 } produces \(\left(\dfrac{i}{n}\right)^{2}\) Just add the words "left" and "right".

OpenStudy (anonymous):

Okay thanks. Do you mind giving mathmale or jim_thompson a medal because I can only give one and honestly they both helped me @tkhunny

OpenStudy (anonymous):

I'll give the other one a medal too

OpenStudy (tkhunny):

No worries.